OpenStudy (ujjwal):

can someone explain me parametric form of a line? My textbook just says parametric form of line is \[\frac{x-x_1}{cos\theta}=\frac{y-y_1}{sin\theta}=r\] and i don't understand it.

OpenStudy (ujjwal):

\(\theta\) is inclination of line.

OpenStudy (anonymous):

since in a right angled triangle we have polar co-ordinates as \[\tan \theta=gradient=\Delta y/\Delta x\] \[\sin \theta/\cos \theta=\Delta y/\Delta x\] cross-multiply \[\Delta x/\cos \theta =\Delta y/\sin \theta\] both these are equal to r since \[\cos \theta=x/r,\sin \theta=y/r\] \[r=\Delta x/\cos \theta=\Delta y/\sin \theta\]

OpenStudy (ujjwal):

So, \(\Delta x\)=\(x-x_1\) for any arbitrary value of x where \(x_1\) is constant. right?

OpenStudy (anonymous):

yes that is correct...

OpenStudy (ujjwal):

thanks mate!

OpenStudy (anonymous):

sure thing